NOTE :

1 mole of a compound is equivalent to its molecular mass

1 mole of a compound is equal to 6.02x10^23 molecules

The percentage is a direct equivalent to grams

Answers should be in 3 significant figures

No need to include the units and compound in the answers


  1. A compound is found to contain 63.52 % iron and 36.48% sulfur. Find its empirical formula.  
  2. A compound is found to contain 26.56% potassium, 35.41% chromium, and the remainder oxygen. Find its empirical formula.  
  3. A 60.00 g sample of tetraethyllead, a gasoline additive, is found to contain 38.43 g lead, 17.83 g carbon, and 3.74 g hydrogen. Find its empirical formula. 
  4. A 170.00 g sample of an unidentified compound contains 29.84 g sodium, 67.49 g chromium, and 72.67 g oxygen. Find the empirical formula.  
  5. Determine the molecular formula of the compound with an empirical formula of CH and a formula mass of 78.110 u.  



1
Expert's answer
2021-06-29T00:39:26-0400

5 .

CH is emperical mass = 12 + 1 = 13 .

n = 78 / 13 = 6 .

Molecular formula = 6 ( CH ) = C6H6 .